# Aligning Large Language Models to Follow Instructions and Hallucinate Less via Effective Data Filtering

### v1

@@ Abstract
Training LLMs on data that contains unfamiliar knowledge during the instruction tuning stage can make LLMs overconfident and encourage hallucinations. To address this challenge, we introduce a novel framework, NOVA, which identifies high-quality data that aligns well with the LLM’s learned knowledge to reduce hallucinations. NOVA includes Internal Consistency Probing (ICP) and Semantic Equivalence Identification (SEI) to measure how familiar the LLM is with instruction data. Specifically, ICP evaluates the LLM’s understanding of the given instruction by calculating the tailored consistency among multiple self-generated responses. SEI further assesses the familiarity of the LLM with the target response by comparing it to the generated responses, using the proposed semantic clustering and well-designed voting strategy. Finally, we introduce an expert-aligned reward model, considering characteristics beyond just familiarity to enhance data quality. By considering data quality and avoiding unfamiliar data, we can utilize the selected data to effectively align LLMs to follow instructions and hallucinate less. Extensive experiments and analysis show that NOVA significantly reduces hallucinations and allows LLMs to maintain a strong ability to follow instructions. [^1] The data and code will be available at https://github.com/S1s-Z/NOVA.
@@
Shuzheng Si♠♢, Haozhe Zhao, Gang Chen, Cheng Gao, Yuzhuo Bai Zhitong Wang, Kaikai An, Kangyang Luo, Chen Qian Fanchao Qi, Baobao Chang, and Maosong Sun Tsinghua University   Peking University   DeepLang AI
@@ 1 Introduction
Certain efforts (Lin et al., 2024b; Zhang et al., 2024b; Tian et al., 2024) apply reinforcement learning (RL) to teach LLMs to hallucinate less after the instruction tuning stage. For example, Zhang et al. (2024b) leverages the self-evaluation capability of an LLM and employs GPT-3.5-turbo (OpenAI, 2022) to create preference data, subsequently aligning the LLM with direct preference optimization (DPO) (Rafailov et al., 2023). However, Lin et al. (2024b) finds that such RL-based methods can weaken the model’s ability to follow instructions. These methods also necessitate additional corpus and API costs of the close-source LLMs, making them inefficient. Different from RL-based methods, an intuitive strategy to align LLMs to follow instructions and hallucinate less is to filter out the instruction data that contains unfamiliar knowledge for the instruction tuning. Unfortunately, previous studies (Liu et al., 2024a; Cao et al., 2024) solely focus on selecting high-quality data to improve the instruction-following abilities of LLMs. Even worse, these selected high-quality data may present more unknown knowledge to the LLM and further encourage hallucinations, as these data may contain responses with expert-level knowledge and often delve into advanced levels of detail.
@@
Therefore, we introduce NOVA, which includes Internal Consistency Probing (ICP) and Semantic Equivalence Identification (SEI), a novel framework designed to identify the instruction samples that align well with LLM’s knowledge, thereby aligning the LLM to follow instructions and hallucinate less. NOVA attempts to measure how well the LLM understands the knowledge in both the given instruction and target response. For ICP, we first prompt the LLM to generate multiple responses to demonstrate what it has learned about a specific instruction during the pre-training stage. Then we use the internal states produced by the LLM to measure how consistent the generated responses are. If the internal states of multiple responses exhibit greater consistency for a given instruction, it indicates that the LLM has internalized the relevant knowledge during pre-training. For SEI, we initially integrate a well-trained model to classify the generated responses that convey the same thing into a semantic cluster. Next, we employ the designed voting strategy to identify which semantic cluster the target response fits in. This helps us find out how many generated responses are semantically equivalent to the target response, indicating how well the LLM understands the target response. If the target response matches well with the largest cluster, it shows the LLM is familiar with its content. Based on ICP and SEI, we can measure how well the model understands the knowledge in instruction data and avoid training it on unfamiliar data, therefore reducing hallucinations. Lastly, we introduce an expert-aligned quality reward model, considering characteristics beyond just familiarity, e.g., the complexity of instructions and the fluency of responses, to enhance data quality. By considering data quality and avoiding unfamiliar data, we can use the selected data to effectively align LLMs to follow instructions and hallucinate less.
@@ 2 Related Work
Data Filtering for Instruction Tuning.  Data are crucial for training neural networks. (Van Engelen and Hoos, 2020; Song et al., 2022; Si et al., 2022, 2023a; Zhao et al., 2024; An et al., 2024; Si et al., 2024a; Cai et al., 2024). According to Zhou et al. (2023), data quality is more important than data quantity in instruction tuning. Thus, many works attempt to select high-quality samples to improve the LLMs’ instruction-following abilities. Chen et al. (2023); Liu et al. (2024a) utilize the feedback from well-aligned close-source LLMs to select samples. Cao et al. (2024); Li et al. (2024a); Ge et al. (2024); Si et al. (2024b); Xia et al. (2024); Zhang et al. (2024a) try to utilize the well-designed metrics (e.g., complexity) based on open-source LLMs to select the samples. However, these high-quality data always contain expert-level responses and may contain much unfamiliar knowledge to the LLM. Unlike focusing on data quality, we attempt to identify the samples that align well with LLM’s knowledge, thereby allowing the LLM to hallucinate less.
@@ 3 Methodology
In this section, we detail our proposed framework NOVA as shown in Figure 2. Previous studies (Lin et al., 2024b; Kang et al., 2024; Gekhman et al., 2024) find that tuning LLMs on data containing new or unfamiliar knowledge can encourage models to be overconfident and lead to hallucinations. Inspired by this finding, NOVA aims to filter out the unfamiliar instruction data for the instruction tuning, thereby aligning the LLM to follow instructions and hallucinate less.
@@ 3.1 Internal Consistency Probing
If the LLM is familiar with the given instruction, the sentence embeddings of generated responses will be highly correlated and the value of $F_{ins}(q)$ will be close to $G$. On the contrary, when the LLM is indecisive, the model will generate multiple responses with different meanings leading to a significant value of $F_{ins}(q)$. In this way, we can exploit the dense semantic information to effectively estimate the knowledge in the given instruction.
@@ 3.2 Semantic Equivalence Identification
Another challenge is to estimate the knowledge in the target response, since the target response can contain expert-level and unfamiliar knowledge for LLMs. Training LLMs on these data can encourage hallucinations. Therefore, we propose Semantic Equivalence Identification (SEI) to measure the LLM’s familiarity with the target response by calculating how many generated responses are semantically equivalent to the target response. If the target response and more generated responses convey the same meaning, it indicates that the LLM is more familiar with it, thereby training the LLM on this target response will reduce hallucinations.
As the target response is manually labeled or derived from advanced LLMs (e.g., GPT-4) instead of generated by the LLM itself, the internal states of the LLM cannot effectively represent the target response. Thus, unlike utilizing internal states as the proposed ICP, we calculate LLM’s familiarity with target responses using the proposed semantic clustering strategy. In detail, we first attempt to cluster the generated responses that convey the same thing into a semantic cluster. This is because instruction data responses are often free-form, and multiple generated responses can have the same meaning in different ways. Therefore, we employ an off-the-shelf natural language inference (NLI) model to cluster these responses. NLI models are trained to infer the logical entailment between an arbitrary pair of sentences. Thus, NLI models are well-suited to identify semantic equivalence, as two generated responses mean the same thing if you can entail (i.e. logically imply) each from the other (Kuhn et al., 2023; Jung et al., 2024). In this way, we can use an NLI model to consider two responses that can be entailed from each other as semantically equivalent responses. Specifically, we test each pair $(r^{\prime}_{i},r^{\prime}_{j})$ of $i$-th and $j$-th generated responses as:
@@
We cluster the generated responses $[r^{\prime}_{1},...,r^{\prime}_{K}]$ into $M$ different semantic clusters $[c_{1},...,c_{M}]$, where $m$-th semantic cluster $c_{m}$ contains $k_{m}$ generated responses. Each semantic cluster $c$ is a set of generated responses that convey the same thing. We further apply the NLI model to determine which semantic cluster the target response $r$ fits in. Specifically, we use the model to test the target response $r$ and each generated response $r^{\prime}_{i}\in[r^{\prime}_{1},...,r^{\prime}_{K}]$:
@@
In this way, we can calculate how many generated responses in a semantic cluster are semantically equivalent to the target response. For semantic clusters $[c_{1},...,c_{M}]$, the counts of such generated responses are $[k^{\prime}_{1},k^{\prime}_{2},...,k^{\prime}_{M}]$. We use the votes in each semantic cluster to decide which cluster the target response belongs to:
@@
We calculate the ratio of the number of responses $k_{target}$ in the target cluster to the total number of generated responses as the final score $F_{res}(r)$:
@@
According to Eq.(8), when the LLM is familiar with the knowledge within the target response $r$, most of the generated responses will have the same meaning as target response $r$, thus the value of $F_{res}(r)$ will be close to $1$. On the contrary, if the target response contains unseen knowledge, i.e., none of the generated responses have the same meaning as it, the value of $F_{res}(r)$ will be close to 00. To this end, we can effectively estimate the knowledge in the target response.
@@ 3.3 Ranking, Selecting, and Training
To comprehensively consider both the knowledge within the instruction and the target response, we estimate the knowledge in the instruction data by calculating the ratio between $F_{ins}(q)$ and $F_{res}(r)$ for an instruction data $(q,r)$:
@@ 4.1 Setup
Baselines.  We compare several strong baselines, including (1) Vanilla Instruction Tuning: Vanilla - 100% fine-tunes the model on the whole instruction dataset; (2) Instruction Data Filtering Methods: IFD (Li et al., 2024a) proposes instruction-following difficulty to select a subset of instruction data. CaR (Ge et al., 2024) further considers both the data quality and diversity by separately introducing the quality-scoring and clustering models. Nuggets (Li et al., 2024b) focuses on selecting high-quality data by identifying samples that notably boost the performance of different tasks after being learned as one-shot instances; (3) RL-based Methods: FLAME-DPOfact (Lin et al., 2024b) introduces atomic fact decomposition and retrieval augmented claim verification to construct preference data and apply DPO. SELF-EVAL (Zhang et al., 2024b) leverages the self-evaluation capability of LLMs and employs GPT-3.5-turbo to create preference data, aligning the LLM with DPO. We apply these RL-based methods after tuning LLMs on the whole instruction dataset.
Implementation Details.  Our main experiments are conducted on LLaMA-3-8B and LLaMA-3-70B (Grattafiori et al., 2024). More details are shown in Appendix B, e.g., the training of quality reward model and hyperparameters.
@@ 4.2 Main Results
NOVA Significantly Reduces Hallucinations.  As shown in Table 1, NOVA shows consistent and significant improvements on three hallucination benchmarks measuring factuality and faithfulness. Compared to indiscriminately using the whole instruction dataset (i.e., Vanilla - 100%), using samples selected by NOVA to train LLMs can improve 3.5-8.6% on BioGEN, 2.0-4.5% on LongFact, and 4.9-8.5% on FollowRAG-Faithfulness. This is because NOVA effectively filters out the unfamiliar instruction data and avoids training LLMs on these data thereby reducing the hallucinations. Compared to instruction data filtering methods that focus on data quality, like IFD, our method consistently improves the performance across different selected sample ratios (5-15%) on three benchmarks. Meanwhile, these data selected by quality-focused methods may present unfamiliar knowledge to the LLM and encourage hallucinations on LongFact. On the contrary, NOVA aims to identify the samples that align well with LLM’s knowledge, helping the LLM to hallucinate less. NOVA also achieves better performance than RL-based methods without introducing additional preference data. These findings underline the effectiveness of our method in aligning LLMs to hallucinate less.
@@ 4.3 Analysis
Variant Methods Testing.  As shown in Table 5, we further explore the variant methods in measuring the LLM’s familiarity. For ICP, we separately replace it with sentence-level confidence (Perplexity) and lexical metrics (Rouge-L). Specifically, we use the average perplexity score generated responses to represent sentence-level confidence and use the average Rouge-L score between each pair of two generated responses as lexical metrics. However, these straightforward strategies neglect highly concentrated semantic information within the internal states, and thus fail to capture the fine-grained differences between responses and limit the final performance. We also explore the effectiveness of an advanced embedding model, we use text-embedding-3-large [^2] https://platform.openai.com/docs/guides/embeddings from OpenAI and set the dimension as 4096. We find that using the internal states achieves better performance, showing the effectiveness of our method. This is because internal states may reflect more dense and fine-grained information from LLM itself that may have been lost in the decoding phase of the responses. For SEI, we explore whether using k-means clustering based on internal states computed as ICP and sentence embedding from text-embedding-3-large can identify suitable semantic clusters. We can find that our method achieves better performance because the k-means algorithm is not based on semantic equivalence to get the clusters. Also, the internal states of LLMs cannot efficiently represent the target response, as this response is manually labeled or generated by other advanced LLMs instead of generated by the LLM itself. We also find that simply voting based on the textual contents instead of semantic clustering limits the final performance, as these responses are often free-form and can have the same meaning in different ways.
Discussion. We conduct the parameter study to test the robustness of our method in Appendix C. We also conduct a transferability study in Appendix D and find NOVA can fit in other LLMs. We further explore the design of our method in Appendix E and find our design is effective.

### v2

@@ Abstract
Training LLMs on data containing unfamiliar knowledge during the instruction tuning stage can encourage hallucinations. To address this challenge, we introduce NOVA, a novel framework designed to identify high-quality data that aligns well with the LLM’s learned knowledge to reduce hallucinations. NOVA includes Internal Consistency Probing (ICP) and Semantic Equivalence Identification (SEI) to measure how familiar the LLM is with instruction data. Specifically, ICP evaluates the LLM’s understanding of the given instruction by calculating the tailored consistency among multiple self-generated responses. SEI further assesses the familiarity of the LLM with the target response by comparing it to the generated responses, using the proposed semantic clustering and well-designed voting strategy. Finally, to ensure the quality of selected samples, we introduce an expert-aligned reward model, considering characteristics beyond just familiarity. By considering data quality and avoiding unfamiliar data, we can utilize the selected data to effectively align LLMs to follow instructions and hallucinate less. Extensive experiments and analysis show that NOVA significantly reduces hallucinations and allows LLMs to maintain a strong ability to follow instructions. [^1] The data and code will be available at https://github.com/S1s-Z/NOVA.
@@
Shuzheng Si♠♢, Haozhe Zhao, Gang Chen, Cheng Gao, Yuzhuo Bai Zhitong Wang, Kaikai An, Kangyang Luo, Chen Qian Fanchao Qi♠♢, Baobao Chang, and Maosong Sun Tsinghua University   Peking University   DeepLang AI
@@ 1 Introduction
Certain efforts (Lin et al., 2024b; Zhang et al., 2024b; Tian et al., 2024) apply reinforcement learning (RL) to teach LLMs to hallucinate less after the instruction tuning stage. For example, Zhang et al. (2024b) leverages the self-evaluation capability of an LLM and employs GPT-3.5-turbo (OpenAI, 2022) to create preference data, subsequently aligning the LLM with direct preference optimization (DPO) (Rafailov et al., 2023). However, Lin et al. (2024b) finds that such RL-based methods can weaken the model’s ability to follow instructions. These methods also necessitate additional preference data and API costs from the advanced LLMs, making them inefficient. Different from RL-based methods, an intuitive strategy to align LLMs to follow instructions and hallucinate less is to filter out the instruction data that contains unfamiliar knowledge for the instruction tuning. Unfortunately, previous studies (Liu et al., 2024a; Cao et al., 2024) solely focus on selecting high-quality data to improve the instruction-following abilities of LLMs. Even worse, these selected high-quality data may present more unknown knowledge to the LLM and further encourage hallucinations, as these data may contain responses with expert-level knowledge and often delve into advanced levels of detail.
@@
Therefore, we introduce NOVA, which includes Internal Consistency Probing (ICP) and Semantic Equivalence Identification (SEI), a framework designed to identify high-quality instruction samples that align well with LLM’s knowledge, thereby aligning the LLM to follow instructions and hallucinate less. NOVA initially uses ICP and SEI to measure how well the LLM understands the knowledge in the given instruction and target response. For ICP, we prompt the LLM to generate multiple responses to demonstrate what it has learned about a specific instruction during pre-training. Then we use the internal states produced by the LLM to assess how consistent the generated responses are. If the internal states of these responses exhibit greater consistency for the instruction, it indicates that the LLM has internalized the relevant knowledge during pre-training. For SEI, we first integrate a well-trained model to classify the generated responses that convey the same thing into a semantic cluster. Next, we employ the designed voting strategy to identify which semantic cluster the target response fits in. This helps us find out how many generated responses are semantically equivalent to the target response, indicating how well the LLM understands the target response. If the target response matches well with the largest cluster, it shows the LLM is familiar with its content. Based on ICP and SEI, we can measure how well the model understands the knowledge in instruction data and avoid training it on unfamiliar data to reduce hallucinations. Lastly, to ensure the quality of selected samples, we introduce an expert-aligned quality reward model, considering characteristics beyond just familiarity, e.g., the complexity of instructions and the fluency of responses. By considering data quality and avoiding unfamiliar data, we can use the selected data to effectively align LLMs to follow instructions and hallucinate less.
@@ 2 Related Work
Data Filtering for Instruction Tuning.  Data are crucial for training neural networks. (Van Engelen and Hoos, 2020; Song et al., 2022; Si et al., 2022, 2023a; Zhao et al., 2024; An et al., 2024; Si et al., 2024a; Cai et al., 2024). According to Zhou et al. (2023), data quality is more important than data quantity in instruction tuning. Therefore, many works attempt to select high-quality instruction samples to improve the LLMs’ instruction-following abilities. Chen et al. (2023); Liu et al. (2024a) utilize the feedback from well-aligned close-source LLMs to select samples. Cao et al. (2024); Li et al. (2024a); Ge et al. (2024); Si et al. (2024b); Xia et al. (2024); Zhang et al. (2024a) try to utilize the well-designed metrics (e.g., complexity) based on open-source LLMs to select the samples. However, these high-quality data always contain expert-level responses and may contain much unfamiliar knowledge to the LLM. Unlike focusing on data quality, we attempt to identify the samples that align well with LLM’s knowledge, thereby allowing the LLM to hallucinate less.
@@ 3 Methodology
In this section, we will detail our proposed framework NOVA as shown in Figure 2. Previous studies (Lin et al., 2024b; Kang et al., 2024; Gekhman et al., 2024) find that tuning LLMs on data containing new or unfamiliar knowledge can encourage models to be overconfident and further lead to hallucinations. Inspired by this finding, NOVA aims to filter out the unfamiliar instruction data for the instruction tuning, thereby aligning the LLM to follow instructions and hallucinate less.
@@ 3.1 Internal Consistency Probing
If the LLM is familiar with the given instruction, the sentence embeddings of generated responses will be highly correlated and the value of $F_{ins}(q)$ will be close to $G$. On the contrary, when the LLM is indecisive, the model will generate multiple responses with different meanings leading to a significant value of $F_{ins}(q)$. In this way, we can exploit the dense semantic information to effectively measure the LLM’s familiarity with the instruction.
@@ 3.2 Semantic Equivalence Identification
Another challenge is to estimate the knowledge in the target response and measure the LLM’s familiarity with it, since the target response can contain expert-level and unfamiliar knowledge for the LLM. Training LLMs on such data can encourage hallucinations. Therefore, we propose Semantic Equivalence Identification (SEI) to measure the LLM’s familiarity with the target response by calculating how many generated responses are semantically equivalent to the target response. If the target response and more generated responses convey the same meaning, it indicates that the LLM is more familiar with it, thereby training the LLM on this target response will reduce hallucinations.
As the target response is manually labeled or derived from advanced LLMs (e.g., GPT-4) instead of generated by the LLM itself, the internal states of the LLM cannot effectively represent the target response. Thus, unlike utilizing internal states as the proposed ICP, we calculate LLM’s familiarity with target responses using the proposed semantic clustering strategy. In detail, we first cluster the generated responses that convey the same thing into a semantic cluster. This is because these responses are often free-form, and multiple generated responses can have the same meaning in different ways. Therefore, we employ an off-the-shelf natural language inference (NLI) model to cluster these responses. NLI models are trained to infer the logical entailment between an arbitrary pair of sentences. Thus, NLI models are well-suited to identify semantic equivalence, as two generated responses mean the same thing if you can entail (i.e. logically imply) each from the other (Kuhn et al., 2023; Jung et al., 2024). In this way, we can use an NLI model to consider two responses that can be entailed from each other as semantically equivalent responses. Specifically, we test each pair $(r^{\prime}_{i},r^{\prime}_{j})$ of $i$-th and $j$-th generated responses as:
@@
In this way, we can identify the semantic equivalence of each pair of generated responses and then cluster these generated responses $[r^{\prime}_{1},...,r^{\prime}_{K}]$ into $M$ different semantic clusters $[c_{1},...,c_{M}]$, where $m$-th semantic cluster $c_{m}$ contains $k_{m}$ generated responses. Each semantic cluster $c$ is a set of generated responses that convey the same thing. We further apply the NLI model to determine which semantic cluster the target response $r$ fits in. Specifically, we use the model to test the target response $r$ and each generated response $r^{\prime}_{i}\in[r^{\prime}_{1},...,r^{\prime}_{K}]$:
@@
Using this method, we can determine how many generated responses in a semantic cluster are semantically equivalent to the target response $r$. For semantic clusters $[c_{1},...,c_{M}]$, the counts of such generated responses are $[k^{\prime}_{1},k^{\prime}_{2},...,k^{\prime}_{M}]$. We use the votes in each semantic cluster to decide which cluster the target response belongs to:
@@
We calculate the ratio of the number of responses $k_{target}$ in the target cluster $c_{target}$ to the total number of generated responses as $F_{res}(r)$:
@@
According to Eq.(8), when the LLM is familiar with the knowledge within the target response $r$, most of the generated responses will have the same meaning as target response $r$, thus the value of $F_{res}(r)$ will be close to 1. On the contrary, if the target response contains unseen knowledge, i.e., none of the generated responses have the same meaning as it, the value of $F_{res}(r)$ will be close to 0. To this end, we can effectively measure the LLM’s familiarity with the target response.
@@ 3.3 Ranking, Selecting, and Training
To comprehensively estimate the knowledge and consider both the LLM’s familiarity with the instruction and the target response, we calculate the ratio between $F_{ins}(q)$ and $F_{res}(r)$ for an instruction data $(q,r)$ as the final score:
@@ 4.1 Setup
Baselines.  We compare several strong baselines, including (1) Vanilla Instruction Tuning: Vanilla - 100% fine-tunes the model on the whole instruction dataset; (2) Instruction Data Filtering Methods: IFD (Li et al., 2024a) proposes instruction-following difficulty to select a subset of instruction data. CaR (Ge et al., 2024) simultaneously considers the data quality and diversity by introducing two scoring methods. Nuggets (Li et al., 2024b) focuses on selecting high-quality data by identifying samples that notably boost the performance of different tasks after being learned as one-shot instances; (3) RL-based Methods: FLAME-DPOfact (Lin et al., 2024b) introduces atomic fact decomposition and retrieval augmented claim verification to construct preference data and apply DPO. SELF-EVAL (Zhang et al., 2024b) leverages the self-evaluation capability of LLMs and employs GPT-3.5 to create preference data, aligning the LLM with DPO. We apply these RL-based methods after tuning LLMs on the whole instruction dataset.
Implementation Details.  Our main experiments are conducted on LLaMA-3-8B and LLaMA-3-70B (Grattafiori et al., 2024). More implementation details are shown in Appendix B, e.g., the training of quality reward model and hyperparameters.
@@ 4.2 Main Results
NOVA Significantly Reduces Hallucinations.  As shown in Table 1, NOVA shows consistent and significant improvements on three hallucination benchmarks measuring factuality and faithfulness. Compared to indiscriminately using the whole instruction dataset (i.e., Vanilla - 100%), using samples selected by NOVA to train LLMs can improve 3.5-8.6% on BioGEN, 2.0-5.1% on LongFact, and 4.9-8.5% on FollowRAG-Faithfulness. This is because NOVA effectively filters out the unfamiliar instruction data and avoids training LLMs on these data thereby reducing the hallucinations. Compared to instruction data filtering methods that focus on data quality, like IFD, our method consistently improves the performance across different selected sample ratios (5-15%) on three benchmarks. Meanwhile, these data selected by quality-focused methods may present unfamiliar knowledge to the LLM and encourage hallucinations on LongFact. On the contrary, NOVA aims to identify the samples that align well with LLM’s knowledge, helping the LLM to hallucinate less. NOVA also achieves better performance than RL-based methods without introducing additional preference data. These findings underline the effectiveness of our method in aligning LLMs to hallucinate less.
@@ 4.3 Analysis
Variant Methods Testing.  As shown in Table 5, we further explore the variant methods in measuring the LLM’s familiarity. For ICP, we separately replace it with sentence-level confidence (Perplexity) and lexical metrics (Rouge-L). Specifically, we use the average perplexity score of generated responses to represent sentence-level confidence and use the average Rouge-L score between each pair of two generated responses as lexical metrics. However, these straightforward strategies neglect highly concentrated semantic information within the internal states, and thus fail to capture the fine-grained differences between responses and limit the final performance. We also explore the effectiveness of an advanced embedding model, we use text-embedding-3-large [^2] https://platform.openai.com/docs/guides/embeddings from OpenAI and set the dimension as 4096. We find that using the internal states achieves better performance, showing the effectiveness of our method. This is because internal states may reflect more dense and fine-grained information from LLM itself that may have been lost in the decoding phase of the responses. For SEI, we explore whether using k-means clustering based on internal states computed as ICP and sentence embedding from text-embedding-3-large can identify suitable semantic clusters. We can find that our method achieves better performance because the k-means algorithm is not based on semantic equivalence to get the clusters. Also, the internal states of LLMs cannot efficiently represent the target response, as this response is manually labeled or generated by other advanced LLMs instead of generated by the LLM itself. We also find that simply voting based on the textual contents instead of semantic clustering limits the final performance, as these responses are often free-form and can have the same meaning in different ways.
Discussion. We conduct the parameter study to test the robustness of our method in Appendix C. We also conduct a transferability study in Appendix D and find NOVA can fit in other LLMs. We further explore the design of our method in Appendix E and find our design is effective. We conduct a case study in Appendix G to qualitatively show the difference between samples with different scores.